# The Universal Theory of Locally Universal Tracial von Neumann Algebras is not Computable

### v1

@@ Abstract
Building on Lin’s breakthrough $\mathsf{MIP}^{\mathsf{co}}=\mathsf{coRE}$ and an encoding of non-local games as universal sentences in the language of tracial von Neumann algebras, we show that locally universal tracial von Neumann algebras have undecidable universal theories. This implies that no such algebra admits a computable presentation. Our results also provide, for the first time, explicit examples of separable $\mathrm{II}_{1}$ factors without computable presentations, and in fact yield a broad family of them, including McDuff factors, factors without property $\Gamma$, and property (T) factors. We also obtain analogous results for locally universal semifinite von Neumann algebras and tracial $C^{*}$-algebras. The latter provides strong evidence for a negative solution to the Kirchberg embedding problem. We discuss how these are obstructions to approximation properties in the class of tracial/semifinite von Neumann algebras.
@@ 1 Introduction
The Connes Embedding Problem (CEP), first posed as an offhand comment in [Con76], asked whether the hyperfinite tracial von Neumann algebra $\mathcal{R}$ is locally universal for tracial von Neumann algebras. Equivalently, does every separable tracial von Neumann algebra embed into an ultrapower of $\mathcal{R}$? At a conceptual level, CEP asked whether all tracial von Neumann algebras can be faithfully approximated by finite-dimensional matrix algebras. Because of its connections to Kirchberg’s QWEP Conjecture, Tsirelson’s Problem in quantum information, the Microstates Conjecture in free probability, and the hyperlinearity problem for groups, CEP stood for decades as one of the central open problems in operator algebras.
@@
In forthcoming work, Lin proves the dual result $\mathsf{MIP}^{\mathsf{co}}=\mathsf{coRE}$ [Lin25]. From an operator algebraic perspective, this is arguably more striking: while $\mathsf{MIP}^{*}=\mathsf{RE}$ yields results about the class of Connes embeddable tracial von Neumann algebras, $\mathsf{MIP}^{\mathsf{co}}=\mathsf{coRE}$ yields results about the class of all tracial von Neumann algebras. In [Man25], the second author suggested that $\mathsf{MIP}^{\mathsf{co}}=\mathsf{coRE}$ should imply that the class of tracial von Neumann algebras admits no uniform approximation property. Our main result formalizes this prediction:
@@
To our knowledge, these are the first explicit examples of separable $\mathrm{II}_{1}$ factors (or even tracial von Neumann algebras) without computable presentations. Such examples have been long sought by computable model theorists. Our methods yield not just a single example, but an entire family: by varying the construction, we obtain explicit $\mathrm{II}_{1}$ factors without computable presentations across natural subclasses, including McDuff factors, factors without property $\Gamma$, and factors with property (T).
@@ Acknowledgements
We would like to thank Isaac Goldbring for providing several helpful comments and corrections on an earlier draft and, in particular, for pointing us to tracial $\mathrm{C}^{*}$-algebras, ultimately leading to Section 4.3. We would also like to thank Junqiao Lin for helpful discussions about $\mathsf{MIP}^{\mathsf{co}}=\mathsf{coRE}$.
@@ 2 Preliminaries
We work in the framework of continuous logic [Yaa+08], specialized to tracial von Neumann algebras (see [FHS14]). See [Gol23] for a detailed discussion of continuous model theory in operator algebras.
@@
  • Elements of the form $p(a_{i_{1}},\ldots,a_{i_{k}})$ where $p$ is a -polynomial with coefficients in $\mathbb{Q}(i)$ and $a_{i_{1}},\ldots,a_{i_{k}}$ are special points are called generated points of the presentation.
  • We say that a presentation $\mathcal{N}^{\#}$ is computable if there is an algorithm which, upon the input of a generated point $p$ of $\mathcal{N}^{\#}$ and a natural number $k$, returns a rational number $q$ such that $|\|p\|_{2}-q|<2^{-k}$.
Intuitively, this means that $\mathcal{N}^{\#}$ generates $\mathcal{N}$ in such a way that it is algorithmically possible to compute the norm of a generated point from a description of how it is generated. See [GH21] for more details on computable presentations of tracial von Neumann algebras.
@@ 3.1 Non-local Games
In this section, we recall the connection between traces on group $C^{*}$-algebras and strategies for non-local games. Our notion of a non-local game is what is commonly called a synchronous game in the literature. Since we will never need non-synchronous games, we will drop the adjective and simply say “non-local game.” For background and a more detailed discussion, see [Gol21, Sections 4–5].
@@
A non-local game $\mathfrak{G}$ consists of:
@@
To analyze non-local games algebraically, consider the universal $C^{*}$-algebra $\mathcal{A}_{Q,A}$ generated by projections
@@
Such a strategy is called a synchronous quantum commuting strategy. We will be dropping the adjective “synchronous” as we do not consider non-synchronous strategies. We can do this by [Lin23, Theorem 1.2] or [MS23, Corollary 3.2]. This definition is equivalent to the standard definition by [Pau+16, Proposition 5.6].
Let $\omega^{\mathrm{co}}(\mathfrak{G})$ denote the best winning probability of $\mathfrak{G}$ over all quantum commuting strategies. In a upcoming paper [Lin25], Lin will prove the following:
@@
  • If $M$ does not halt, then $\omega^{\mathrm{co}}(\mathfrak{G}_{M})=1$;
  • If $M$ halts, then $\omega^{\mathrm{co}}(\mathfrak{G}_{M})\leq\tfrac{1}{2}$.
@@
Note again that above we are able to use the synchronous value by [Lin23, theorem 1.2] or [MS23, Corollary 3.2].
@@ 3.2 Universal Sentences for Game Values
We now explain how to convert the quantum commuting value of a non-local game into the maximum value of a universal sentence in the language of tracial von Neumann algebras. This section follows a lot of ideas from [GH24, Section 3].
@@
Given a non-local game $\mathfrak{G}=(Q,A,\mu,D)$ with $|Q|=n$ and $|A|=m$, and a tuple $\bar{e}=(e_{x}^{a})_{x\in Q,\ a\in A}\in X_{n,m}^{\mathcal{N}}$ inside a tracial von Neumann algebra $(\mathcal{N},\tau)$, define
@@
For a non-local game $\mathfrak{G}$, the commuting-operator value satisfies
$\omega^{\mathrm{co}}(\mathfrak{G})\;=\;\sup_{(\mathcal{N},\tau)}\ \sup_{\bar{e}\in X_{n,m}^{\mathcal{N}}}\ \psi_{\mathfrak{G}}(\bar{e}),$
@@
($\geq$) Suppose $\bar{e}\in X_{n,m}^{\mathcal{N}}$ is a PVM in a tracial von Neumann algebra $(\mathcal{N},\tau)$. By universality, there is a *-homomorphism $\pi:\mathcal{A}_{Q,A}\to\mathcal{N}$ sending the generating PVM in $\mathcal{A}_{Q,A}$ to $\bar{e}$. Pulling back the trace, $\tau\circ\pi$ is a tracial state on $\mathcal{A}_{Q,A}$. The corresponding correlation $p(a,b|x,y)=\tau(e_{x}^{a}e_{y}^{b})$ is therefore a quantum commuting strategy, and its value is precisely $\psi_{\mathfrak{G}}(\bar{e})$. This shows that every $\bar{e}$ contributes to $\omega^{\mathrm{co}}(\mathfrak{G})$, hence
$\omega^{\mathrm{co}}(\mathfrak{G})\;\geq\;\sup_{(\mathcal{N},\tau)}\ \sup_{\bar{e}\in X_{n,m}^{\mathcal{N}}}\ \psi_{\mathfrak{G}}(\bar{e}).$
($\leq$) Conversely, let $p$ be any quantum commuting strategy for $\mathfrak{G}$. By definition, there exists a tracial state $\tau$ on $\mathcal{A}_{Q,A}$ such that $p(a,b|x,y)=\tau(e_{x}^{a}e_{y}^{b})$. Since the trace simplex of $\mathcal{A}_{Q,A}$ is compact convex, the value $\omega^{\mathrm{co}}(\mathfrak{G})$ is attained at some trace. The GNS of this is a tracial von Neumann algebra and the generating PVM of $\mathcal{A}_{Q,A}$ gets sent to a PVM in this algebra. Thus
$\omega^{\mathrm{co}}(\mathfrak{G})\;\leq\;\sup_{(\mathcal{N},\tau)}\ \sup_{\bar{e}\in X_{n,m}^{\mathcal{N}}}\ \psi_{\mathfrak{G}}(\bar{e})$
@@
Fix a Turing machine $M$. By [Lin25], there is a non-local game $\mathfrak{G}_{M}$ computable from $M$ such that $\omega^{\mathrm{co}}(\mathfrak{G}_{M})=1$ if $M$ does not halt, and $\omega^{\mathrm{co}}(\mathfrak{G}_{M})\leq\tfrac{1}{2}$ if $M$ halts.
@@
$\sup_{\mathcal{N}}\varphi_{M}^{\mathcal{N}}\;=\;\omega^{\mathrm{co}}(\mathfrak{G}_{M}).$
@@ 4.2 Results on Computability of Universal Theories
The above theorem provides, to our knowledge, the first explicit examples of separable $\mathrm{II}_{1}$ factors (or even tracial von Neumann algebras) with no computable presentations. It is easily seen via cardinality considerations that the vast majority of separable $\mathrm{II}_{1}$ factors do not admit computable presentations. More precisely, since there are only countably many computable presentations and far more than countably many separable $\mathrm{II}_{1}$ factors, most separable $\mathrm{II}_{1}$ factors will not admit computable presentations. There are known examples of $\mathrm{II}_{1}$ factors whose so-called “standard presentations” are not computable, but this does not rule out the possibility that they admit some other presentation which is computable.
Recall from [FHS14a] that given any locally universal factor $\mathcal{S}$, both $\mathcal{S}\otimes\mathcal{R}$ and $\mathcal{S}*\mathcal{R}$ are also locally universal. The former is McDuff and the latter does not even have property $\Gamma$. We can also embed $\mathcal{S}$ into a property (T) factor by [CDI23]. Thus we have also demonstrated the first explicit examples of McDuff $\mathrm{II}_{1}$ factors, $\mathrm{II}_{1}$ factors without property $\Gamma$, and $\mathrm{II}_{1}$ factors with property (T) that do not admit computable presentation.
This may provide a post-hoc explanation for why operator algebraists have, to this point, been unable to find a satisfactory representation of a non-Connes embeddable $\mathrm{II}_{1}$ factor. That is, the most obvious candidates (the locally universal $\mathrm{II}_{1}$ factors) cannot be presented in a way that is computable.
@@ 4.3 Consequences for $\mathrm{C}^{*}$-algebras
A natural question is to what extent our methods extend to the setting of $\mathrm{C}^{*}$-algebras. Since our proofs rely on traces, we work in the language of tracial $\mathrm{C}^{*}$-algebras. We need a $\mathrm{C}^{*}$-algebra analogue of Lemma 3.3 which, in turn, depends on the stability results of either [KPS18] or [Sal22]. Neither of these stability results immediately generalize because they both rely on spectral projections, which are not guaranteed to exist in $\mathrm{C}^{*}$-algebras.
On the other hand, an encouraging result in this direction is [ESS20, Theorem 3.7], which implies definability of the set of PVMs in each $\mathrm{C}^{*}$-algebra since the PVM $\mathrm{C}^{*}$-algebra is $C^{*}(\mathcal{F}_{Q,A})$, and $\mathcal{F}_{Q,A}$ is virtually free. However, this definability is not uniform and the result does not give an explicit almost-near modulus, let alone a computable one.
Our next lemma, together with the discussion to follow, directly constructs the necessary almost-near modulus.
@@
$\omega^{\mathrm{co}}(\mathfrak{G})\ =\ \sup_{(\mathcal{A},\tau)}\ \sup_{\overline{e}\in X_{n,m}^{\mathcal{A}}}\ \psi_{\mathfrak{G}}\bigl{(}\overline{e}\bigr{)}.$
@@
We are thus tantalizingly close to a negative solution of the Kirchberg embedding problem. However, achieving such a full resolution using our methods would require removing the dependence on traces. This seems unlikely, as the operator norm alone carries too little information about the value of a quantum strategy. An entirely new approach appears to be necessary.

### v2

@@ Abstract
Building on Lin’s breakthrough $\mathsf{MIP}^{\mathsf{co}}=\mathsf{coRE}$ and an encoding of non-local games as universal sentences in the language of tracial von Neumann algebras, we show that locally universal tracial von Neumann algebras have undecidable universal theories. This implies that no such algebra admits a computable presentation. Our results also provide, for the first time, examples of separable $\mathrm{II}_{1}$ factors without computable presentations, and in fact yield a broad family of them, including McDuff factors, factors with neither property $\Gamma$ nor property (T), and property (T) factors. We also obtain analogous results for locally universal semifinite von Neumann algebras and tracial $C^{*}$-algebras. The latter provides strong evidence for a negative solution to the Kirchberg embedding problem. We discuss how these are obstructions to approximation properties in the class of tracial/semifinite von Neumann algebras.
@@ 1 Introduction
The Connes Embedding Problem (CEP), first posed as an offhand comment in [Con76], asked whether the hyperfinite tracial von Neumann algebra $\mathcal{R}$ is locally universal for tracial von Neumann algebras. That is, does every separable tracial von Neumann algebra embed into an ultrapower of $\mathcal{R}$? At a conceptual level, CEP asked whether all tracial von Neumann algebras can be faithfully approximated by finite-dimensional matrix algebras. Because of its connections to Kirchberg’s QWEP Conjecture, Tsirelson’s Problem in quantum information, the Microstates Conjecture in free probability, and the hyperlinearity problem for groups, CEP stood for decades as one of the central open problems in operator algebras.
@@
Lin recently announced a proof of the dual result $\mathsf{MIP}^{\mathsf{co}}=\mathsf{coRE}$ in a preprint [Lin25]. From an operator algebraic perspective, this is arguably more striking: while $\mathsf{MIP}^{*}=\mathsf{RE}$ yields results about the class of Connes embeddable tracial von Neumann algebras, $\mathsf{MIP}^{\mathsf{co}}=\mathsf{coRE}$ yields results about the class of all tracial von Neumann algebras. In [Man25], the second author suggested that $\mathsf{MIP}^{\mathsf{co}}=\mathsf{coRE}$ should imply that the class of tracial von Neumann algebras admits no uniform approximation property. Our main result is a model theoretic formulation of this prediction:
@@
Suppose $\mathsf{MIP}^{\mathsf{co}}=\mathsf{coRE}$, then for each Turing machine $M$, there is a restricted universal sentence $\sigma_{M}$ in the language of tracial von Neumann algebras, computable from the description of $M$, such that:
@@
The above theorem provides, to our knowledge, the first examples of separable $\mathrm{II}_{1}$ factors (or even tracial von Neumann algebras) with no computable presentations. It is easily seen via cardinality considerations that the vast majority of separable $\mathrm{II}_{1}$ factors do not admit computable presentations. There are known examples of $\mathrm{II}_{1}$ factors whose so-called “standard presentations” are not computable, but this does not rule out the possibility that they admit some other presentation which is computable. Our methods yield not just a single example, but an entire family: by varying the construction, we obtain $\mathrm{II}_{1}$ factors without computable presentations across natural subclasses, including McDuff factors, factors with neither property $\Gamma$ nor property (T), and factors with property (T).
@@ Acknowledgements
We would like to thank Isaac Goldbring for providing several helpful comments and corrections on an earlier draft and, in particular, for pointing us to tracial $\mathrm{C}^{*}$-algebras, ultimately leading to Section 4.3. We would also like to thank Junqiao Lin for helpful discussions about $\mathsf{MIP}^{\mathsf{co}}=\mathsf{coRE}$. Finally, we thank the anonymous referees for helpful comments.
@@ 2 Preliminaries
We work in the framework of continuous logic [Ben+08], specialized to tracial von Neumann algebras (see [FHS14]). See [Gol23] for a detailed discussion of continuous model theory in operator algebras.
@@
  • Elements of the form $p(a_{i_{1}},\ldots,a_{i_{k}})$ where $p$ is a -polynomial with coefficients in $\mathbb{Q}(i)$ and $a_{i_{1}},\ldots,a_{i_{k}}$ are special points are called rational points of the presentation.
  • We say that a presentation $\mathcal{N}^{\#}$ is computable if there is an algorithm which, upon the input of a rational point $p$ of $\mathcal{N}^{\#}$ and a natural number $k$, returns a rational number $q$ such that $|\|p\|_{2}-q|<2^{-k}$.
Intuitively, this means that $\mathcal{N}^{\#}$ generates $\mathcal{N}$ in such a way that it is algorithmically possible to compute the norm of a rational point from a description of how it is generated. See [GH21] for more details on computable presentations of tracial von Neumann algebras.
@@ 3.1 Non-local Games
In this section, we recall the connection between traces on group $C^{*}$-algebras and strategies for non-local games. For background and a more detailed discussion, see [Gol21, Sections 4–5].
@@
A synchronous non-local game $\mathfrak{G}$ consists of:
@@
To analyze synchronous non-local games algebraically, consider the universal $C^{*}$-algebra $\mathcal{A}_{Q,A}$ generated by projections
@@
Such a strategy is called a synchronous quantum commuting strategy. This definition is equivalent to the standard definition by [Pau+16, Proposition 5.6].
Let $\omega^{s}_{\mathrm{co}}(\mathfrak{G})$ denote the best winning probability of $\mathfrak{G}$ over all synchronous quantum commuting strategies. In [Lin25], Lin announced a proof of the following:
@@
  • If $M$ does not halt, then $\omega^{s}_{\mathrm{co}}(\mathfrak{G}_{M})=1$;
  • If $M$ halts, then $\omega^{s}_{\mathrm{co}}(\mathfrak{G}_{M})\leq\tfrac{1}{2}$.
@@
Note that above we are able to use the synchronous value by [Lin23, theorem 1.2] or [MS23, Corollary 3.2].
@@ 3.2 Universal Sentences for Game Values
We now explain how to convert the synchronous quantum commuting value of a synchronous non-local game into the maximum value of a universal sentence in the language of tracial von Neumann algebras. This section follows a lot of ideas from [GH24, Section 3].
@@
Given a synchronous non-local game $\mathfrak{G}=(Q,A,\mu,D)$ with $|Q|=n$ and $|A|=m$, and a tuple $\bar{e}=(e_{x}^{a})_{x\in Q,\ a\in A}\in X_{n,m}^{\mathcal{N}}$ inside a tracial von Neumann algebra $(\mathcal{N},\tau)$, define
@@
For a synchronous non-local game $\mathfrak{G}$, the commuting-operator value satisfies
$\omega^{s}_{\mathrm{co}}(\mathfrak{G})\;=\;\sup_{(\mathcal{N},\tau)}\ \sup_{\bar{e}\in X_{n,m}^{\mathcal{N}}}\ \psi_{\mathfrak{G}}(\bar{e}),$
@@
($\geq$) Suppose $\bar{e}\in X_{n,m}^{\mathcal{N}}$ is a PVM in a tracial von Neumann algebra $(\mathcal{N},\tau)$. By universality, there is a *-homomorphism $\pi:\mathcal{A}_{Q,A}\to\mathcal{N}$ sending the generating PVM in $\mathcal{A}_{Q,A}$ to $\bar{e}$. Pulling back the trace, $\tau\circ\pi$ is a tracial state on $\mathcal{A}_{Q,A}$. The corresponding correlation $p(a,b|x,y)=\tau(e_{x}^{a}e_{y}^{b})$ is therefore a synchronous quantum commuting strategy, and its value is precisely $\psi_{\mathfrak{G}}(\bar{e})$. This shows that every $\bar{e}$ contributes to $\omega^{s}_{\mathrm{co}}(\mathfrak{G})$, hence
$\omega^{s}_{\mathrm{co}}(\mathfrak{G})\;\geq\;\sup_{(\mathcal{N},\tau)}\ \sup_{\bar{e}\in X_{n,m}^{\mathcal{N}}}\ \psi_{\mathfrak{G}}(\bar{e}).$
($\leq$) Conversely, let $p$ be any synchronous quantum commuting strategy for $\mathfrak{G}$. By definition, there exists a tracial state $\tau$ on $\mathcal{A}_{Q,A}$ such that $p(a,b|x,y)=\tau(e_{x}^{a}e_{y}^{b})$. Since the trace simplex of $\mathcal{A}_{Q,A}$ is compact convex, the value $\omega^{s}_{\mathrm{co}}(\mathfrak{G})$ is attained at some trace. The GNS of this is a tracial von Neumann algebra and the generating PVM of $\mathcal{A}_{Q,A}$ gets sent to a PVM in this algebra. Thus
$\omega^{s}_{\mathrm{co}}(\mathfrak{G})\;\leq\;\sup_{(\mathcal{N},\tau)}\ \sup_{\bar{e}\in X_{n,m}^{\mathcal{N}}}\ \psi_{\mathfrak{G}}(\bar{e})$
@@
Fix a Turing machine $M$. By [Lin25], there is a synchronous non-local game $\mathfrak{G}_{M}$ computable from $M$ such that $\omega^{s}_{\mathrm{co}}(\mathfrak{G}_{M})=1$ if $M$ does not halt, and $\omega^{s}_{\mathrm{co}}(\mathfrak{G}_{M})\leq\tfrac{1}{2}$ if $M$ halts.
@@
$\sup_{\mathcal{N}}\varphi_{M}^{\mathcal{N}}\;=\;\omega^{s}_{\mathrm{co}}(\mathfrak{G}_{M}).$
@@ 4.2 Results on Computability of Universal Theories
The above theorem provides, to our knowledge, the first examples of separable $\mathrm{II}_{1}$ factors (or even tracial von Neumann algebras) with no computable presentations. It is easily seen via cardinality considerations that the vast majority of separable $\mathrm{II}_{1}$ factors do not admit computable presentations. More precisely, since there are only countably many computable presentations and far more than countably many separable $\mathrm{II}_{1}$ factors, most separable $\mathrm{II}_{1}$ factors will not admit computable presentations. There are known examples of $\mathrm{II}_{1}$ factors whose so-called “standard presentations” are not computable, but this does not rule out the possibility that they admit some other presentation which is computable.
Recall from [FHS14a] that given any locally universal factor $\mathcal{S}$, both $\mathcal{S}\otimes\mathcal{R}$ and $\mathcal{S}*\mathcal{R}$ are also locally universal. The former is McDuff and the latter has neither property $\Gamma$ nor property (T). We can also embed $\mathcal{S}$ into a property (T) factor by [CDI23]. Thus we have also demonstrated the first examples of McDuff $\mathrm{II}_{1}$ factors, $\mathrm{II}_{1}$ factors with neither property $\Gamma$ nor property (T), and $\mathrm{II}_{1}$ factors with property (T) that do not admit computable presentation.
Our results may explain in retrospect why the poor-man’s solution to the Connes Embedding Problem in [FHS14] has yet to yield direct operator-algebraic constructions of non-Connes-embeddable factors. Our results say that the examples constructed via compactness cannot easily be used in computations. We remark on the other hand that there are many factors that are conjecturally not Connes-embeddable but nevertheless have computable presentations. For example, the group von Neumann algebra of the Higman group is expected to not be Connes-embeddable (cf. [Tho12]) but admits a computable presentation by the decidability of the word problem for the Higman group [DLU12] and [GH21]. Even in this case, our results rule out a potential proof strategy for showing it is not Connes-embeddable: the group von Neumann algebra of the Higman group cannot be locally universal.
@@ 4.3 Consequences for $\mathrm{C}^{*}$-algebras
A natural question is to what extent our methods extend to the setting of $\mathrm{C}^{*}$-algebras. Since our proofs rely on traces, we work in the language of tracial $\mathrm{C}^{*}$-algebras. We need a $\mathrm{C}^{*}$-algebra analogue of Lemma 3.3 which, in turn, depends on the results of either [KPS18] or [Sal22] who each use spectral projections to establish 2-norm stability of POVMs.
Fortunately, this problem essentially reduces to a simple unitary stability result for $\mathrm{C}^{*}$-algebras. Here, the norm control of polar decompositions is the correct $\mathrm{C}^{*}$-analog of spectral projection arguments in 2-norm.
@@
$\omega^{s}_{\mathrm{co}}(\mathfrak{G})\ =\ \sup_{(\mathcal{A},\tau)}\ \sup_{\overline{e}\in X_{n,m}^{\mathcal{A}}}\ \psi_{\mathfrak{G}}\bigl(\overline{e}\bigr).$
@@
We are thus tantalizingly close to a negative solution of the Kirchberg embedding problem. However, achieving such a full resolution using our methods would require removing the dependence on traces. An entirely new approach appears to be necessary.